OpenStudy (anonymous):

Determine the solution set of x2 - 80 = 0

OpenStudy (anonymous):

do i solve this by saying 80*80

OpenStudy (mimi_x3):

is it \(x^2 - 80 =0\)

OpenStudy (anonymous):

yes

OpenStudy (mimi_x3):

do you konw that you're solving for \(x\)?

OpenStudy (anonymous):

I Have to find the square root

OpenStudy (mimi_x3):

no the \(x\) lets go slowly

OpenStudy (mimi_x3):

\(x^2 - 80 = 0\) \(x^2 = 80\)

OpenStudy (mimi_x3):

do yo understand? when you move it to the other side its a positive?

OpenStudy (anonymous):

yes i understand

OpenStudy (mimi_x3):

then when its squared you remove it to the other side its sqrt

OpenStudy (mimi_x3):

i mean move**

OpenStudy (anonymous):

so i do 80 times 80

OpenStudy (mimi_x3):

\[x^2 = 80\] show me what you would to solve for x

OpenStudy (anonymous):

im not sure if i square 80 or find the factors of 80

OpenStudy (mimi_x3):

why would you square it? its sqrt(80) when you move it to the left hand side

OpenStudy (anonymous):

what is sqrt

OpenStudy (mimi_x3):

square root

OpenStudy (anonymous):

so do i find the factors of it then

OpenStudy (mimi_x3):

why would you want tol find find factors you are just solving for \(x\)!!!

OpenStudy (anonymous):

o well im jus not understanding thats why im asking becasue i don't know my next steps

OpenStudy (mimi_x3):

show me what you did after x^2 =80 in this question you are basically solving for a varaible i.e. x

OpenStudy (anonymous):

i dont know thats why im here to get this explained to me so i can do it by myself

OpenStudy (mimi_x3):

okay i just told you!! the right hand side is x^2 then the left hand side is squareroot

OpenStudy (anonymous):

thank you